OpenStudy (anonymous):

3x^2-576=0 find all real solutions

OpenStudy (anonymous):

3x^2 = 576 x^2 = 192 x = 13.85 (approx)

OpenStudy (anonymous):

I think for this one all you have to use is the quadratic equation

OpenStudy (anonymous):

I have to have 2 answers

OpenStudy (anonymous):

\[-b \pm \sqrt{b^2-4ac} /2a\]

OpenStudy (anonymous):

yes, so that means you must use the equation i gave you to get those to answers

OpenStudy (anonymous):

The two answers will be +13.85 and -13.85.

OpenStudy (anonymous):

Which numbers do I use for which letters?

OpenStudy (anonymous):

\[ax^2+bx+c=0\]

OpenStudy (anonymous):

So looking at the equation you need to solve for match the numbers with the letters. if there is no number there for a letter count it as 0

OpenStudy (anonymous):

Thanks !